OpenStudy (gibbs):

-6 – = -8 whats the intiger?

OpenStudy (anonymous):

-8 + (-6) =?

OpenStudy (gibbs):

-14?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

You can check it by -6 + (-14) = -8

OpenStudy (gibbs):

it was 2

OpenStudy (anonymous):

How.

OpenStudy (anonymous):

-6 - (-14) = -8

OpenStudy (slickster256):

my bad i didnt mean to type that lol